### v1

# Signal Processing over Time-Varying Graphs: A Systematic Review

Abstract
As irregularly structured data representations, graphs have received a large amount of attention in recent years and have been widely applied to various real-world scenarios such as social, traffic, and energy settings. Compared to non-graph algorithms, numerous graph-based methodologies benefited from the strong power of graphs for representing high-dimensional and non-Euclidean data. In the field of Graph Signal Processing (GSP), analogies of classical signal processing concepts, such as shifting, convolution, filtering, and transformations are developed. However, many GSP techniques usually postulate the graph is static in both signal and typology. This assumption hinders the effectiveness of GSP methodologies as the assumption ignores the time-varying properties in numerous real-world systems. For example, in the traffic network, the signal on each node varies over time and contains underlying temporal correlation and patterns worthy of analysis. To tackle this challenge, more and more work are being done recently to investigate the processing of time-varying graph signals. They cope with time-varying challenges from three main directions: 1) graph time-spectral filtering, 2) multi-variate time-series forecasting, and 3) spatiotemporal graph data mining by neural networks, where non-negligible progress has been achieved. Despite the success of signal processing and learning over time-varying graphs, there is no survey to compare and conclude the current methodology for GSP and graph learning. To compensate for this, in this paper, we aim to review the development and recent progress on signal processing and learning over time-varying graphs, and compare their advantages and disadvantages from both the methodological and experimental side, to outline the challenges and potential research directions for future research.
Index Terms:
Graph Signal Processing, Time-Varying Graphs, Graph Machine Learning, Graph Neural Networks
I Introduction
Classical signal processing techniques are developed for signals defined on regular domains, such as time and space, the underlying structure of which can be represented by a uniform grid. In recent years, the processing of signals residing on non-Euclidean domains has gradually become more and more important in real-world practice, where the edges describe the relationships between the signal elements. Motivated by this, Graph Signal Processing (GSP) has arisen as a powerful framework for analyzing and processing signals over irregular domains. The application of GSP has covered many fields including transportation networks [1], temperature networks [2], and brain connectivity networks [3].
Conventional multi-variate algorithms can be applied to signals and data on graphs, as these signals also fall into the category of multi-variate signals. It is often more advantageous to use GSP-based methods when there is a topological structure associated with the data, especially in cases where the graph structure plays a crucial role in the dynamics of the underlying data. Non-graph methods often rely on independent or identically distributed assumptions, which fail to capture the relational dependencies between data points that are naturally encoded in a graph. This urges the development of algorithms and models that are designed specifically to address these limitations of applying conventional methods to graph signals.
Analogous to classical discrete signal processing, GSP extends concepts such as graph shift, graph convolution, Graph Fourier Transform (GFT), Graph Wavelet Transform (GWT), and basic filters like low-pass, high-pass, and band-pass filters. The notions of FIR and IIR filters are also extended in GSP [4, 5, 6, 7]. GSP-based algorithms are capable of solving various classical machine learning tasks, such as classification [8] and clustering [9]. GSP techniques also form the foundation of spectral graph-based deep learning algorithms, such as ChebNet [10] and Graph Convolutional Networks (GCN) [8], where nonlinear activations are combined with GSP methods to enable nonlinear modeling. The GSP components within graph neural networks provide model interpretability, which was previously lacking in non-graphical deep learning algorithms [11].
Nonetheless, the world is dynamic, which makes the signal vary over time, leading to the demand for time-varying signal processing. The signals captured in the real world are often incomplete and corrupted by noise. Let us assume that we are in a challenging missing-not-at-random scenario, where certain nodes consistently have missing data. In practice, this type of missing observation often corresponds to sensors that have failed or been removed from [12]. This scenario makes it difficult for non-graph methods to produce predictions accurately, as inferring model parameters can be challenging due to the missing observations and noise. Furthermore, these consistent missing observations also severely affect the accuracy if the attempt is to use correlation-based methods to infer the signals. Conversely, GSP can leverage the graph’s topological information to conduct spatial inference in addition to the temporal predictions. As an extension of GSP, Time-Varying Graph Signal Processing (TV-GSP) aims to analyze and process the signals that are residing on graphs and evolving with time. This extension is critical in many real-world situations. Examples include traffic flow in urban networks, the spread of information in social networks, and brain activity patterns measured over time. TV-GSP provides the means for understanding how these signals evolve and how they can be leveraged to extract meaningful information. A key challenge in TV-GSP is how to effectively model the joint variation of signals over both the spatial domain and temporal domain. One approach is to treat time as an additional dimension and to extend traditional GSP techniques to the spatiotemporal domain. This involves constructing joint spatiotemporal representations, where the graph structure and temporal dynamics are simultaneously taken into account.
In this review, we aim to provide an overview to present the recent progress in TV-GSP. Specifically, we first classify the current TV-GSP works into three categories according to their methodology. Then, we discuss their advantages and disadvantages in different settings. Finally, we propose several new potential research directions in this area. The main contributions are summarized as follows:
We provide a comprehensive review of the fundamental concepts and characteristics of TV-GSP, construct the knowledge system and systematically investigate the key techniques and models.
We present in detail the current mainstream TV-GSP algorithms and analyze in depth their advantages and disadvantages.
We discuss the challenges faced in TV-GSP highlighting potential future research directions and technological trends.
The remainder of this paper is organized as follows. In Section II, we begin with important definitions and background knowledge about TV-GSP. Sections III, IV, V, and VI separately provide a comprehensive review and discussion of the existing TV-GSP approaches following the taxonomy demonstrated in Figure 1. Section VII summarizes the commonly used datasets, the corresponding application, and evaluation metrics. Finally, Section VIII concludes and indicates the future direction of the TV-GSP.
II Preliminaries
The advantage of using graphs to represent multi-variate signals is that graphs capture and represent the irregular interactions in data. The underlying topological structures of a graph signal are often formed to represent the pairwise relationship between two nodes through the connection of edges. Since an edge between two nodes is a pairwise relationship by definition, we can intuitively form and remove edges based on some measures of the strength of the relationship between the two data on each node. When this process is repeated on all the data pairs, a graph is formed on the global level.
II-A Definition of Time-Varying Graphs and Signals
In conventional GSP that operates on static data, the main research focus is placed on a static graph that is defined by a set of nodes and similarly a set of edges that represent the connections between nodes. In this review, the graph is always assumed as undirected unless otherwise specified, and can be either weighted or unweighted. A graph signal is defined as a multi-dimensional vector, recording the value of each node in . We further denote the number of nodes with .
The adjacency matrix of comprises the topological information by storing the state of connectivity between nodes. To be specific, the entry of is the edge weight between nodes and . If is unweighted, indicates there is an edge between nodes and . Conversely, means there is no connection between and . For an undirected graph, the adjacency matrix is always symmetric. If is undirected and unweighted, the number of edges connected to a node is defined as the node degree , which can be represented by a diagonal matrix called the degree matrix . In the weighted case, the degree of a node is the sum of all edge weights connected to it, rather than simply counting the number of edges.
Since the real-world graph is rarely static and mostly dynamic, where the node values, edge weights, and the topology are all probably evolving over time, applying GSP techniques on every single time instance is obviously improper due to neglect of temporal dependency. Thus, the TV-GSP came into being. However, different TV-GSP algorithms are proposed for different sorts of time-varying graphs. In this review, we additionally divide the time-varying graphs into two classes from the evolutionary perspective. The first class is referred to as the Node-Varying Graphs (NVG), where the node signals vary over time while other graph elements, such as the topology or edge weights, remain static. Both time-vertex graphs, commonly used in GSP, and spatiotemporal graphs, frequently seen in GNNs, can be categorized under NVGs, as they focus on temporal variations in node signals. An illustration of NVG is shown in Fig. 2(a).
The second class is named Dynamic Graph (DG), where the varying signal not only appears at nodes but also happens in the graph topology itself. Correspondingly, the approach for processing NVG is named NVGSP, and the approach for processing dynamic graphs is named DGSP. The representative works in these two classes are shown in Fig. 1. Given multi-dimensional series data , where is the number of vertices and is the number of time points. For such data, time-varying graphs are efficient structures to characterize the relations between different vertices in a certain spatial and temporal range. A time-varying graph can be represented by the notation , where is the vertices set, and respectively denote the edges set and adjacency matrix at time t. In most time-varying graph modeling scenarios, the size of is constant, the size of can be time-varying or time-constant, and also changes with . In general, from the perspective of connectivity, graphs can be directed, undirected, weighted, and unweighted. In this review, we additionally divide the time-varying graphs into two classes from the evolutionary perspective.
II-B Graph Signal Processing
The graph Laplacian matrix , which combines information from and , is defined as . The Graph Fourier Transform (GFT) is based on the eigenvector decomposition
where is the orthonormal eigenvector matrix and is the diagonal matrix of eigenvalues . The GFT transforms a graph signal from the spatial domain to the spectral domain by projecting onto :
A graph signal transformed to the spectral domain can be converted back to the spatial domain using the inverse graph Fourier transform (IGFT):
The GFT of the time-varying graph signal (hourly temperature) in Fig. 2(a) is shown in Fig. 2(b). Spectral-domain operations can be performed similarly to those in the classical Fourier transform by defining a filter , which is then applied using the convolution property of the Fourier transform. A basic yet complete graph convolution procedure to apply the filter to in the spectral domain and transform it back into the spatial domain:
In GSP, concepts like low-pass, high-pass, band-pass, and band-reject filters can be analogously implemented through filter design [7]. Wavelets on graphs can be constructed by applying a kernel function to the decomposition of the graph Laplacian in (1) to achieve localized representations [6]. A graph signal is spectrally sparse when it is bandlimited in the frequency domain. For example, a bandlimiting filter is defined based on a frequency set , in which , with if and 0 otherwise [13].
The spectral graph convolution can be approximated into a spatial graph convolution by using Chebyshev polynomial approximation for a series of Chebyshev polynomial defined over the interval [14]:
with the approximation being
where is the number of polynomials, and are the weight of the polynomial.
A graph signal with a reduced number of nodes, sampled or masked according to a node sampling set , is sparse in the spatial domain. The matrix represents the sampling or masking matrix, containing only non-zero diagonal elements given by . The matrix can be designed to introduce spatial sparsity based on a sampling strategy or it can be used as a masking matrix to record the observation state of the nodes.
| Notation | Definition |
| Underlying graph | |
| The set of nodes | |
| The set of edges | |
| N | The number of nodes |
| F | The number of features in each node |
| The vector of graph signals (node values) | |
| Identical matrix | |
| Adjacency matrix | |
| Degree matrix | |
| Laplacian matrix | |
| Normalized adjacency matrix | |
| Loss function |
III Graph Acquisition and Representation Techniques
What differentiates GSP from conventional signal processing or standard multivariate algorithms is its integration of graph structures and topological relationships alongside signal data. This distinct approach leverages the underlying graph topology to enhance data representation and analysis in GSP. Consequently, the formulation of an appropriate graph topology when time-varying graph signals are present is crucial for representing data within GSP and graph machine learning frameworks. Before delving into the algorithmic aspects of TV-GSP, it is essential to examine key techniques for representing multivariate signals on topological structures.
III-A Conventional Static Graph Acquisition and Representations
The simplest and most common way of acquiring the graph topology is when the signals naturally contain or form a graph-like or network-like structure. Real-world data represented on the nodes can be constructed based on realistic relationships. A common example of naturally occurring graph structures can be found in traffic forecasting tasks: in the real world sensors positioned at various locations along highways that record data such as traffic speed or flow are perceived as nodes in a graph; the edges in the graph represent the connections between nodes, corresponding to the spatial distribution of sensors across the highways [15]. In the case of meteorological data collection, variables such as longitude and latitude from several weather stations are aggregated and analyzed using a Gaussian Kernel to construct a k-nearest-neighbor graph [2]. A graph depicting the strength of 5G signal reception is created as detailed in [13] by documenting signal strength across various sensors. A slightly more abstract category of graph generated from the real world is the social network. For example, users on social media are represented as nodes, and interactions between users create edges between nodes [16]. Another example is a citation network in which each node is a scientific paper and edges represent citations from one paper to another [17]. While many real-world datasets inherently exhibit graph structures, in other cases, the graph topology may not appear as straightforward. Specialized methods are employed to construct graphs that capture underlying relationships in the data.
Under the scenario where direct inference of the graph topology is not straightforward, a graph can still be formed using some techniques. In statistical graphical models, statistical measures between two data elements are used to define the pairwise relationship. For example, the covariance matrix and the precision matrix can be used to form a graph because covariance is a statistical measure of the deviations between two variables, meaning that measuring and properly thresholding this pairwise relationship measure gives us a sparse graph [18]. The Gaussian Graphical Model is a specific example of a Statistical Graphical Model that represents a conditional dependency structure in a graph: the nodes of the graph are the multivariate Gaussian distributed variables, and edges between these variables are the conditional dependencies or non-zero partial correlation coefficients [19, 20]. In GSP operations such as graph construction, transforms, sampling, prediction, and regularization can be realized from a probabilistic and statistical aspect using the Gaussian Graphical Model [21]. The graph Laplacian can be learned using a merging of the Gaussian prior and a smoothness assumption on the multivariate signal, which is equivalent to learning the underlying topological structure [11]. A multivariate signal can be assumed to be a sparse linear combination of atoms from a dictionary based on a polynomial of graph Laplacians that represents localized patterns from a sparse prior [22]. Another sparse prior assumption topology learning approach can be found in [23]: the Gaussian Graphical Model is used in a way that data is trained in a neural network to learn a function calculating the empirical covariance matrices to form graph structures.
Graph embedding techniques leverage adjacency and topological properties to project signals and features onto an embedding space to facilitate the analysis, visualization, and application in downstream GSP tasks. In the spectral domain, graph embeddings can be achieved by leveraging the eigenvectors of graph representations such as the adjacency matrix [24]. In the spatial domain, early approaches to generating the graph embedding are to use the random walk to propagate through the graph and then embed the data on the walk trajectories as words using approaches such as work2vec [25]. The DeepWalk generates embeddings by performing random walks on the graph, treating each node sequence like a sentence in natural language processing [26]. The node2vec extended DeepWalk further by introducing a flexible search mechanism that balances breadth-first and depth-first search strategies [27]. Recently, a more common method to generate graph embeddings is through GNNs such as GCNs. The spectral GCN is achieved by feeding the spectral graph convolution in (4) into a nonlinear activation function and replacing the filter with learnable filters as if the filters are parameters leaned from backpropagation [28]. The spatial GCN and the ChebNet are the spatial domain versions of GNNs with the spectral convolutions approximated by the Chebyshev polynomials as shown in (6) and let the neural network learn the polynomial coefficient instead of the spectral filter [8, 10]. The GraphSAGE exploits graph aggregations that sample and combine features from node neighborhoods, enabling an inductive framework where the model generalizes better to unseen nodes during inference [16]. The message-passing neural networks (MPNNs) further generalize the previous GNNs by introducing a flexible and localized method that computations are defined at the node level based on purely the node local adjacencies instead of defining all computations at once [29]. These graph embedding techniques establish a foundation for integrating GSP with machine learning tasks, including classification and clustering, by transforming graph signals along with the information in the graph topology into representations suitable for non-graph machine learning models.
III-B Dynamic Graph Acquisition and Representation
The pairwise relationships in graphs can be dynamic. Time-evolving processes over graphs can be captured by kernels formed by a preselected dictionary that fits the signal of interest where partial correlation, partial Granger causality, non-linear structural equations, and vector autoregressions techniques are used along with assumptions of the graph signals such as low-rank, sparsity, and smoothness [30, 31]. A dynamic graph can be formed from the correlated properties, which are previously seen in static Statistical Graphical Models, in time-varying graph signals by considering the space-time interactions in the underlying signal [32, 33, 34]. The topology inference on dynamic signals can be approached using a state-space formulation of a network process [35]. The Graphical LASSO is another instance of the Statistical Graphical Model method that estimates the inverse covariance matrix (precision matrix) with sparsity constraints enforced through an penalty. By combining Graphical LASSO with an ADMM-based algorithm, one can incorporate signals from neighboring time points and pairwise connections to learn the smoothly changing graphs [36]. Another smoothness-based graph topology learning task is to use the Sobolev smoothness as a measure to define a new smoothness function for time-varying graph signals [37]. The work of Natali et al. proposes to use the structural equation model to learn graph topologies in an online fashion and further proposes a combination of the Gaussian graphical model, the structural equation model, and the smoothness-based model into an all-in-one online topology learning framework [38, 39]. The causal relationships between data can be modeled using a Vector Autoregressive model, making it possible to infer a graph topology from the underlying time-series vector [40]. A specific example of this approach can be seen in a time-varying gene activation graph formed by thresholding a regression matrix of a time-varying gene regression that reflects gene activation patterns over time, leading to a parametric representation of the time-evolution and influence between pairs of data [41]. Given an underlying graph structure at a one-time instance, assuming that the dynamics of a time-varying graph only have some minor changes between time instances, a small perturbation analysis of the Laplacian matrix and its GFT can be done to model the time-evolution of the dynamic graph [42]. Topology inference of a dynamic topology from time-varying data can be tackled from another angle on an iterative prediction-correction approach on a Gaussian Graphical Model [43].
III-C Representing High-order Signals on Graphs
Despite the success of GSP, data representation in GSP is solely on the graph nodes, meaning that higher-order signals, such as signals on graph edges or signals on structures formed by edges, are not being represented. Examples of signals existing on the graph but not only the graph nodes can be found in the following literature: traffic data of the road network being recorded on the edges [44], gene interaction over time forms a set of time-varying graph edges [41], multi-agent systems that use time-varying edge data to model the agent interactions [45].
Higher-order structures such as hypergraphs, simplicial complexes, cell complexes, and multi-graphs enable the possibilities of defining graph representations beyond pairwise relationships [46]. Researchers have introduced the concept of Topological Signal Processing (TSP) as an alternative perspective to GSP [47]. In TSP, the Hodge Laplacian, a more generalized form of the graph Laplacian, is used to represent higher-order features beyond individual graph nodes [48]. The Hodge Laplacian can be used to represent simplicial complexes or cell complexes [47, 49]. When using simplicial complexes as building blocks, signals on graph nodes are treated as signals on the order simplex, graph edges are treated as signals on the order simplex, and triangles formed by edges are treated as signals on the order simplex, enabling the representation of signals on edges and beyond [47, 50]. Signal smoothing and denoising can be achieved on the graph edges by the simplicial complex representation [51]. Simplicial signals can also be embedded into simplicial complexes using methods similar to graph (node) embedding [52, 53]. Semi-supervised learning is applied to edge data using the Hodge Laplacians and flow conservation principles [44]. An analogous concept of spectral wavelets for simplicial complexes, termed Hodgelets, is introduced in [54]. Topological Slepians were proposed as signals concentrated on the topological domain and localized in the frequency domain [55]. Finite impulse response filters on simplicial complexes are discussed and further defined within convolution operations for simplicial complex signals [56, 57, 58]. A Chebyshev polynomial approximation for filters on simplicial complexes is also presented in [58] and [59]. These advancements underscore the important role of high-order representation in enhancing GSP for complex data.
IV Graph Signal Processing for Time-varying Graphs
When dealing with time-varying graph signals, GSP algorithms can be broadly categorized into online and offline methods. Online GSP algorithms, such as adaptive filters and state-space models, update dynamically with time-varying data, offering real-time processing capabilities. Offline methods, on the other hand, analyze graph signals in batches, leveraging the entire dataset to capture spatiotemporal relationships. In this section, we will focus on non-GNN approaches, covering algorithm genres such as adaptive filtering, time-series models, and their extensions for graph-structured data.
IV-A Online GSP Algorithms
Let us begin with the discussion of online GSP algorithms for time-varying graph signals. The online GSP algorithms share a workflow similar to that of the online algorithms found in non-graph contexts. Following the convention seen in most adaptive GSP algorithms, we assume that the data sequence arrives one instance at a time, with each instance will be denoted as . In some cases, the data can be partially observed, in which the partial observation can be modeled by applying a sampling matrix to the receive time instance . The data received can also contain noise, which is often represented as an additive noise term . Typically, there will be a weight term defined by utilizing the graph and serves as how the observation is modified or processed. Without the loss of generality, we denote this weight term as but keep in mind that the type and dimension of this term are algorithm-specific. In many cases, the weight is updated based on calculating a metric between the estimation and the observation . The online GSP algorithm update under noisy and missing data observations can be generalized using the following framework:
where denotes the graph signal update function and denotes the weight update function. An illustration of the workflow of online GSP algorithms is shown in Figure 3.
1) GSP-based (Spectral) Adaptive filters: A natural idea for the online processing of time-varying node signals is to utilize adaptive graph filters that consist of a combination of classical adaptive filters with graph shift operations and can efficiently process time-varying node signals owing to their simplicity of implementation. Similar to the classical adaptive filters, GSP-based adaptive filters update in the direction opposite to the error at each time step based on convex optimization. What makes this combination different from the classical adaptive filters is that most GSP-based adaptive filters utilize a predefined bandlimited filter in the spectral domain derived from GFT, which implies that it captures the topological information of the graph in the spectral domain along with the time-varying signal. Using the framework we defined in (7) and the graph convolution in (4), the weight in GSP-based adaptive filters can be expressed as . By leveraging topological information, GSP-based adaptive filters enable signal reconstruction across both spatial and temporal dimensions. The first GSP-based adaptive filter is the adaptive graph Least Mean Squares (GLMS) algorithm[13]. In the GLMS, the update strategy is the optimization results derived based on a -norm optimization with the assumption of Gaussian noise:
where is the expectation, and is the static predefined bandlimited filter. As a general rule of thumb, the update function of the GSP-based adaptive filter can be derived using gradient methods:
The plugging (8) into (9), update function of GLMS can be expressed as
where is the step-size parameter. There are several variants and extensions of the GLMS. The GLMS was originally proposed on a spatiotemporal graph where the graph signals are time-varying but the topology is static; by introducing small perturbations to the graph Laplacian matrix, the GLMS can be extended to dynamic graphs [60]. A variant of the GLMS is the Graph Normalized LMS (GNLMS), in which a spectral domain normalization is derived from minimizing the spectral error between consecutive estimations and then introduced to the weights of the LMS-based update function to improve the convergence behavior [2]. Another variant is the graph recursive least squares (GRLS), which leverages the RLS framework to provide enhanced convergence behavior and superior prediction performance compared to GLMS [61, 62]. By recursively minimizing a weighted least squares error criterion, GRLS adapts more quickly and more accurately to time-varying changes under noisy signals when compared to GLMS.
Gaussian noise assumption is seen in most noise models, and -norm optimization is the go-to option for Gaussian noise because minimizing the squared error corresponds to the maximum likelihood estimate solution [63]. However, the underlying noise in a variety of realistic applications, including meteorological recordings [64] and powerline communication [65], is verified to possess impulsive behaviors that could be represented by heavy-tailed, non-Gaussian distributions, such as generalized Gaussian, Student’s t, and -stable distributions [63]. The least-squares-based methods lose their validity in impulsive noise situations owing to the presence of large or infinite variance caused by the impulsiveness in the noise [66]. Several approaches can be taken to overcome this drawback. Instead of using second power, a least mean power is used to define a loss based on the assumption that the noise is symmetric -stable distribution (SS), giving us the Graph Least Mean algorithm [67]. The Normalized GLMP was further proposed to include a time-varying spectral normalization matrix to speed up the computation of GLMP [68]. Moving further with the idea of -norm optimization, the adaptive Graph-Sign algorithm (Graph-Sign) was derived as a graph extension of the classical adaptive sign error or the least mean absolute deviation (LMAD) algorithm for multivariate signals [69, 70]. The Graph-Sign algorithm is derived from the minimum dispersion criterion and then reduced to -norm optimization, which removes the need for prior knowledge from any noise assumption [71]. This allows the Graph-Sign algorithm to avoid the instability seen in the least-squares-based algorithms when estimating graph signals under impulsive noise.
with and being the regularization term. Taking GLMP as a specific example, the GLMP update function following the framework shown in (7) can be obtained by plugging (11) into (9):
where is the Hadamard product and the superscript p-1 denotes element wise power.
An alternative approach than simply modifying the loss function by switching -norm to -norm or -nor to enhance robustness is to employ in which a robust adaptive estimation of graph signals is developed by replacing the conventional of using -norm to define optimization problem with the Welsch loss [72]. The Welsch loss differs from conventional loss functions by incorporating robustness as a continuous parameter, allowing the Welsch loss to adapt dynamically during algorithm deployment [73]. Instead of directly changing the cost function, a less extreme method to the problem is to add a regularization term to the LMS cost function: the Graph Alternating Minimization algorithm improves the robustness of GLMS by separately modeling the impact of Gaussian noise and outliers in the cost function [74]. The above GSP-based adaptive filters can be further generalized to have the cost function
2) GSP-based (Spatial) Adaptive filters: Notice that the adaptive GSP algorithms are not limited to the spectral domain but could also be conducted in the spatial domain by using the Chebyshev polynomial approximation shown in (6). For example, the bandlimited filters in the GLMS can be approximated using a series of Chebyshev polynomials, leading to the spatial domain graph diffusion [75]. A GSP-based adaptive filter that uses a combination of graph diffusion and sign update strategy was recently proposed in [76]. Recently, the Graph Signal Adaptive Message Passing has been proposed, which is an adaptive filter defined using the flexible localized node message passing scheme instead of defining all computations globally on the graph level [77]. In general, we can use the Chebyshev polynomial approximation shown in (6) to define the weight in the cost function such as the ones seen in (8) and (11):
where are the coefficients for after merging the terms within . It is worth mentioning that the spatial methods resemble the structure of many GNNs, as the majority of GNNs are derived based on the (spatial) GCN [8, 10]. In short, one layer of GCN can be obtained by some algebraic manipulations of (6) and then feeding its output into an activation function. We will be conducting a more thorough review of GNN-based methods for time-varying graph signals in Section V.
3) State Space Models for TV-GSP: Aside from GSP-based adaptive filters, another direction to approach the problem is to use state space models, and the typical example is the Kalman Filter [78]. The Kalman Filter is a recursive algorithm used to estimate the state of a linear dynamic system under the Gaussian noise assumption. In GSP, we can formulate system equations suitable for time-varying graph signals by slightly modifying the framework in (7):
where represents the state evolution function, and represents the observation function. Both functions operate on graph signals using the Laplacian matrix . In the system equations (14), the terms and are independent white Gaussian noises with covariance matrices and , respectively.
Following the approach commonly used in GSP-adaptive filters, the graph structure can be fully leveraged by applying the GFT in (2), which transforms the system equations in (14) into the spectral domain [79]:
The operations within and can now be conducted entirely in the spectral domain using spectral graph convolution in (4). With the multi-variate signals represented as graph signals, the rest of the procedure, such as the update step and the prediction step, follows the classical Kalman filter workflow, with adjustments for the graph representation. Once the estimation and update steps are completed, the results are transformed back into the spatial domain using the Inverse Graph Fourier Transform (IGFT) in (3).
The Kalman Filter over signals on sensor networks can be established for which the optimal estimation method at each sensor is done when the sensor communications are formed by an undirected graph [80]. The Extended Kalman filter (EKF) extends the applicability of the standard Kalman filter from linear systems to non-linear systems by locally approximating them as linear [81]. The graph version of the Extended Kalman filter is a graph process obeying a nonlinear state-space model that can be found by forming a GSP-based Bayesian filter whose goal is to minimize the computed covariance matrix as the Kalman gain of the Extended Kalman filter in the graph spectral domain [79]. The Unscented Kalman Filter (UKF) addresses the non-linearity of systems by representing the mean and covariance of the state using a set of deterministically chosen sample points to avoid the need for linearization, which is different from the standard Kalman filter [82]. This approach can also be applied to graph signals, which leads to the Unscented Kalman filter of graph signals by again using the decomposition of graph Laplacian matrix to form the GFT and then design the Kalman gain matrix from the eigenvectors [83]. However, the Kalman filter is designed for Gaussian noise because of minimizes the mean square error of the estimated states, which is not suitable for non-Gaussian noise.
Remarks Online algorithms are most preferred when the outputs are required to be made on the fly as new inputs are received. Their ability to produce real-time outputs based on very limited data observations in the temporal domain, without requiring the entire input upfront, makes them highly effective in TV-GSP tasks due to the fact that most time-varying graph signals already have a large spatial dimension. Online GSP algorithms are not limited to regression tasks; their outputs can also serve as inputs for classification or clustering models [31, 84]. This capability broadens the application scope of online GSP algorithms. On one hand, online algorithms often have lower computational complexity to cope with real-time output capability, which is a critical advantage of deploying online algorithms on time-varying graph signals. On the other hand, this same online requirement often restricts the length of the prediction horizon of online algorithms, limiting their ability to generate long-term temporal predictions. In contrast, offline algorithms process the entire dataset at once, enabling long-term predictions on graph signals. In the next subsection, we will be discussing offline GSP-based algorithms for time-varying graph signals.
IV-B Offline GSP Algorithms
1) General Offline GSP Methods: One naive way to process the time-varying signals is to apply the graph convolution in (4) or a GCN to the observations as we receive new observations of the graph signal. However, this method does not capture any temporal changes in the signal, which often results in poor performance. To effectively process time-varying signals on graphs, it is essential to capture both spatial and temporal dynamics.
Another way to use GSP to process time-varying graph signals is to treat each time instance as a feature and formulate a multi-feature GSP cost function to minimize the error between estimation and ground truth :
where is the Frobenius Norm, is the regularization weight, Tr is the trace, and is a spatial-smoothness-based regularization that minimizes the total variation of the results [5]. The solution can be found by calculating the gradient and setting it to zero. This method has 2 potential drawbacks. First, treating signals from different time instances as features means that the method is still inherently time-invariant, which does not take into account any temporal dynamics within the signal. Second, after some algebraic manipulations, the final solution requires the calculation of the matrix inverse, which can be unstable and unobtainable when the number of nodes in a graph is very large. This first limitation can potentially mitigated by replacing the spatial smoothness in (16) with the Sobolev Smoothness, which captures the smoothness in both the spatial domain and the temporal domain [37]. To address the second limitation, one can approximate the matrix inverse using Newton’s method, which gives us an iterative method that can be implemented in a distributed manner [85].
2) Joint Time and Graph Fourier Transform: The foundational work on graph wavelets [6] laid the groundwork for localized spectral analysis on static graph signals, which later evolved into more comprehensive spatiotemporal frameworks. To handle scenarios where graph signals evolve over time, it is essential to consider both spatial and temporal dimensions simultaneously using GSP approaches. This has led to the development of Time-vertex methods and the Joint Time and Graph Fourier Transform (JFT) that integrate time-domain information directly into the spectral operations seen in the GSP framework [86]. The Joint Time-Vertex Fourier Transform captures both the spatial domain variations and the temporal domain variations by combining GFT and DFT, providing a comprehensive analysis of spatiotemporal signals, merging classical signal processing and GSP into a unified framework [87, 88]. The JFT also made the definition of dynamic wavelets on graphs possible for tracking the evolving time-varying graph signals [89]. This allows it to effectively decompose signals into joint time-vertex frequency components, offering insights into both temporal dynamics and graph-structured relationships. To ensure stability and enhance performance, an additional regularization term is introduced, promoting smoothness across both graph and time dimensions [88].
3) Time-series Analysis on Graphs: As we explained earlier, conventional multivariate time series models, such as Vector Autoregressive (VAR) and Vector Autoregressive Moving Average (VARMA) processes, can be applied directly to graph signals since these signals are inherently multivariate [90]. However, conventional multivariate time series models are primarily designed for temporal tasks, which limits their ability to exploit the spatial relationships captured within the graph topologies. Taking the VARMA model as an example, it has the expression of
where is the AR order, is the MA order, is a multi-variate random variable, is the by AR parameter, and is the by MA parameter, is an identity matrix, and denotes the matrix of white noise at time . It becomes difficult to accurately infer model parameters in (17) when there are nodes consistently missing, causing conventional VARMA models to underperform. To tackle temporal variability and represent space-time cross-dimensional interactions, one can merge time series analysis methods with GSP.
Time-series analysis algorithms were introduced into GSP by specifying their parameters in the spectral domain via GFT, using the multivariate time-series as graph signals and potentially decreasing parameter size compared to traditional VAR and VARMA [91, 92]. The GSP analogy of VAR and VARMA is the graph Vector Autoregressive (GVAR) model [91] and the graph Vector Autoregressive–Moving-Average (GVARMA) model [93] respectively. Specifically, in VAR, the temporal association is modeled by an AR process, and the spatial association is captured by describing the matrix coefficients of the AR process as graph polynomial filters. In GVARMA, the VARMA coefficients are designed independently from the underlying graph but instead are as spectral filters. The filter design now takes into account concurrently both the spatial and temporal domains:
In (18), to incorporate the spatial information through GSP, the parameters are now defined using graph convolution (4):
Additionally, these parameters can also be defined using spatial graph methods by using the Chebyshev approximation (6) [92]:
Parameter reduction in GVARMA can be achieved as follows. Even though both VARMA and GVARMA parameters have size by , the GVARMA in (19) defines the parameters in the spectral domain as shown in (19) by using the filter , which has only parameters. The number of parameters can be further reduced by defining a frequency set with , and then defining parameters only for the frequencies that present in . As a result, the actual size of the parameters that the GVARMA needs to learn in (19) is drastically reduced to . Similarly, if the GVARMA was to be derived from the spatial domain as shown in (20), the actual sizes of the parameters that the GVARMA needs to learn are and in and respectively, for all and .
VARMA and GVARMA models are limited by their assumption of constant covariance, which makes them insufficient for capturing time-varying features such as volatility. Volatility, a key concept in finance, has been extensively studied by researchers. In time series analysis, models such as Autoregressive Conditional Heteroscedasticity (ARCH) [94] and its generalized version (GARCH) [95] were proposed to address symmetric volatility. To capture asymmetric volatility, the exponential GARCH (EGARCH) model [96] and the GJR-GARCH model [97] were subsequently developed. Recently, a unified framework for defining GARCH models on graphs was introduced [98]. The G-GARCH model shares the same model formula as GVARMA in (18) with the exception that the covariance matrix of also follows a GVARMA model instead of having constant variance. The G-GARCH model treats graph signals and their residuals as graph-stationary processes [99]. By leveraging the graph spectral domain, G-GARCH decomposes a multivariate GARCH model into a linear combination of several univariate GARCH processes, enabling efficient processing of graph-structured data. The G-GARCH model can also be derived into asymmetric models such as the G-EGARCH and G-GJR-GARCH [98].
V Graph Neural Networks for Time-Varying Graphs
Time-varying graphs can be effectively processed using signal-processing-based techniques, which offer powerful tools for analyzing and understanding graph-structured data. These GSP methods can be further enhanced by addressing certain challenges. They usually rely on priors for the parameters and filter design for simple and efficient implementation, but the priors are sometimes inaccessible. Additionally, augmenting traditional linear GSP techniques with mechanisms to capture complex non-linear patterns can significantly improve their ability to handle network data with intricate dependencies and dynamics.
With the recent revival of Neural Networks (NN) [100, 101, 102], Graph Convolutional Networks (GCN) was proposed, which combined nonlinear activation functions and learnable parameters with graph convolution operators in (6) and attained promising performance for many tasks over graphs, such as node classification, link prediction, and node regression [8, 10, 28, 103]. Despite the success of GCN in processing general static graph data, when the graph varies over time, the GCN cannot read the temporal information and fails to produce satisfying predictions. Therefore, developing graph neural networks that can adaptively learn and process time-varying graphs has been an increasingly attended issue. In this chapter, we summarize the existing works for TV-GSP and classify them into three technical directions: spatiotemporal neural network architectures, dynamic neural network architectures, and graph continual learning.
V-A Spatiotemporal Graph Neural Networks
1) Decoupled Learning Methods: Learning spatial and temporal features respectively is the most natural idea for dealing with time-varying challenges. Based on this idea, several decoupled spatiotemporal graph neural networks are proposed [15, 104, 105]. The spatiotemporal Graph Convolutional Networks (STGCN) [15] is the most typical decoupled learning method, and it mainly comprises two parts for the sake of successive spatiotemporal learning, which can be expressed as:
where and separately represent the temporal and graph convolutional operators. They are defined as:
where and are two feature maps generated from by two 1-D convolution operators (i.e., one for feature representation and one for gating). can be the feature maps of the last neural network layer; represents the -th learnable parameter, and applies the -th order Chebyshev polynomial [10] to the graph Laplacian matrix.
2) Integrated Learning Methods: Compared with successive learning spatial and temporal features by two independent modules, merging spatial and temporal feature learning into one layer can be more conducive to learning joint information. Motivated by this, a series of integrated spatiotemporal graph neural networks are proposed [106, 107, 108], where the GCN is usually employed as a spatial feature extractor in an RNN layer for joint spatiotemporal learning. The Diffusion Convolutional Recurrent Neural Networks (DC-RNN) proposed to insert a tailored diffusion convolution into each RNN layer [106]. The proposed diffusion convolution can be modeled as:
where and are two independent parameters for the transition matrices of diffusion process and reverse one respectively. Then, the diffusion convolution with activation function is used to replace the matrix multiplications in the Gated Recurrent Units (GRU) [109] to construct a Diffusion Convolution GRU (DCGRU) as a basic unit in its sequential model. By this means, the temporal information and spatial information is processed collaboratively.
3) Noise-Aware Methods: Some methods also take into account the noise-induced data missing in modeling, and the representative work is the spatiotemporal Graph Gated Recurrent Units (SG-GRU) [110] that are developed to make prediction temporally and complete the missing nodes simultaneously with the aid of GFT and GRU. To achieve this, it first calculates the graph interpolation operator with a given admissible graph sampling operator :
where means the minimal singular value of the given matrix; is the sub-matrix of the eigenvector matrix with columns restricted to the indices associated with the frequencies indicated in ; is assumed as bandlimited.
By this constraint, the graph interpolation operator can be found and used to recover the corrupted graph signal after forecasting as:
where and are predicted by GRU and Spectral GRU from respectively. The major drawback of this empirical approach is that the GFT only acts as a domain transformation and the performance relies only on the GRU, which is not adapted to the graphs and lacks interpretability. Besides, the hyper-parameters selection (e.g., , , and ) heavily depends on the prior knowledge of the graph.
Remarks: Constructing a Spaito-Temporal Graph Neural Network (STGNN) to collaboratively learn and process the time-varying features is a direct way to cope with time-varying challenges. As shown in Fig. 5(a), the paradigm of STGNNs can be summarized as consisting of a pair of time modules (e.g., 1D convolution, RNN and GRU) and spatial modules (e.g., GCN, ChebNet, and GAT), and an extra spatiotemporal fusion module. The difference between various STGNNs generally appears in the architecture design of these modules and how spatial and temporal features are fused. Some methods also consider denoising and inpainting in the process of spatiotemporal graph learning [111].
V-B Dynamic Graph Neural Networks
Most of the STGNNs are designed for node-level time-varying signal processing, while the temporal variation on higher-order signals (e.g., edges and topologies) are overlooked. Recently, Dynamic Graph Neural Networks (DGNN) were proposed to address this problem. In DGNNs, the time-varying graph is divided into two classes. One is the Discrete-Time Dynamic Graph (DTDG), and another is the Continuous-Time Dynamic Graph (CTDG). In DTDG, the dynamic graph is represented by a series of snapshots that can be the graph adjacency matrices and corresponding node feature vectors over time. In CTDG, the time-varying graph is represented by an initial graph and a sequence of graph evolution events (e.g., node insertion, edge addition, and edge deletion). By different dynamic graph modeling, various DGNNs are introduced. We then discuss them separately.
1) DTDG Methods: As illustrated by Fig. 5(b), the typical pipeline of DGNNs for processing DTDG is drawn upon the experience from the architecture of RNN and can be described as:
where represents the current state of timestamp , which can be learned from the current graph topology (i.e., the adjacency matrix); represents the hidden state at timestamp . and are two distinct neural networks. In general, is a GCN, and can be a linear projection layer. Compared to the STGNN that assumes fixed graph topology, in DTDG, the evolving graph topology is considered as variables to assist learning at every timestamp to accurately reflect the node relationship with time [112, 113].
Jointly learning the graph topology and node embedding spatiotemporally at each timestamp is sometimes awkward and not scalable when dealing with large-scale graphs. To relieve this issue, considering the advantage of the Auto-Encoder (AE) in terms of information compression [114], AE is also involved in processing DTDG. Particularly, DynGEM [115] generates the graph embedding sequentially and incrementally based on the encoded information from the last timestamp to enhance scalability and grant the model with inherent ability to capture temporal correlation between adjacent snapshots. On the other hand, dyngraph2vec [116] attempts to encode the graph snapshots within a short time window as a whole based on the temporal smoothness assumption to improve efficiency.
2) CTDG Methods: Since the graph representation is totally different from DTDG, where the event set stores the temporal information rather than iteratively forwarding previous information like RNN. As a consequence, the DGNN methods for CTDG are more diverse. A class of methods models the event sequence by temporal point process, where the occurrence of events recording the evolving graph signal is assumed to follow a multivariate temporal point process. The conditional intensity function of this process can be modulated by the score of the relevant relationship, and the measurement of the relationship can be realized by GNNs [117, 118]. Besides the temporal point process, as an extension of random walk [26, 27], the temporal random walk is utilized for learning spatiotemporal association as well. In temporal random walk, the movement from one vertex to another will depend on its spatiotemporal neighbors to extract dynamics from the time-varying graph [119, 120].
On the other hand, some researchers also model the relationship between temporal events directly by neural networks. For instance, the Dynamic Graph Neural Networks (DyGNN) [121] employ an update component and a propagation component to cope with coming events. As shown in Fig. 5(c), when a new event comes, the update component maintains the freshness of node information by learning the relationship with previously memorized events and the time interval between the stored events and the new event. Then, the propagation component propagates the updated information to the affected nodes (e.g., its neighborhood nodes) by considering the intensity of the influence. Meanwhile, JODIE [122] is developed and shares a similar scheme with DyGNN, which also has the update operation and the projection operation. The major difference is, in JODIE, each node has a static embedding to represent its fixed attributes, and a dynamic embedding to reflect its current state.
Remarks: DGNN is a sort of more flexible NN method for learning and processing time-varying graphs compared with STGNNs, by considering the potential temporal variation in high-order structures of time-varying graphs. However, more complicated modeling generally leads to higher complexity. For instance, in the traffic network, the topology is hard to change, whereas STGNNs are more popular owing to their efficiency. Thus, selecting a proper class of models according to the graph data encountered in practice is critical.
V-C Graph Lifelong Learning
Graph lifelong learning [123] (a.k.a., continual graph learning [124]) is an emerging research topic and mainly developed for processing real-world time-growing graphs. The scene of graph lifelong learning is that the number of nodes is growing over time, or there will be new nodes of unseen classes. As depicted in Fig. 5(c), it solves this problem by iteratively training the GNN. Like general continual learning, the core problem in graph lifelong learning is also catastrophic forgetting. To prevent this, the current graph lifelong learning methods mostly refer to continual learning and can be classified into three categories including the architectural methods, the regularization methods, and the experience-replay methods.
1) Architectural Methods: The architectural methods [123, 125] aim to change the architecture of the GNN iteratively to coherently adapt to the changes that occurred in streaming data. The Feature Graph Networks (FGN) [123] are a special type of architectural method to tackle time-growing graphs by node-feature transformation. Specifically, it first transforms each node into a feature graph, and the features in this node become new nodes in each feature graph. The edge weights between these feature nodes are determined by cross-correlation. Specifically, if the feature node matrix is , then the feature adjacency matrix can be defined as:
where represents the k-hop neighbor of , indexes the channel. In this way, the continually coming nodes can be regarded as an individual graph and classified with previous techniques in continual learning.
2) Experience-Replay Methods: In addition to architectural methods, like in continual learning, the experience replay technique is also employed for graph continual learning [126, 127]. In the Experience Replay Graph Neural Networks (ER-GNN) [126], a set of nodes are stored in a buffer as experience and replayed in the subsequent task to avoid forgetting. To identify valuable nodes of each task, ER-GNN proposes several metrics including the mean of features, coverage maximization, and influence maximization. Once the nodes are selected and stored in a buffer, the training objective for each task is altered as:
where indicates the loss function (e.g., cross entropy in node classification tasks); indexes the task; represents the experience buffer; denotes the graph neural network; represents the training dataset of task .
3) Regularization Methods: Adding regularization terms to penalize dramatic updates on important parameters is also common in continual learning [124, 128, 129]. The representative one in this line of work is, a plug-and-play module, called “Topology-aware Weight Preserving” (TWP) [124], which estimates an importance score for each network parameter. This score is computed as the sum of gradients concerning parameters (denoted as ) and gradients concerning attention coefficients parameters (denoted as ). The first term aims to minimize loss of different tasks, and it can be calculated from:
where is the training feature embedding of task ; is the network parameter; is the loss function. For the second term, it is calculated from:
where represents a single neural network layer for feature projection; represents the concatenated feature embedding of node and at -th layer in the graph attention neural network.
Remarks: Dealing with the time-varying graph by graph continual learning is a noteworthy direction, which merges techniques from both continual learning and graph neural networks and achieves great performance in some specific scenarios (e.g., learning and processing the time-growing graphs [123]).
| Type | Dataset | Domain | Catagory | #Nodes | #Edges | #Timestamps |
| STG | METR-LA [106] | Traffic | general | 207 | 1,515 | 34,272 |
| PeMSBay [106] | Traffic | general | 325 | 2,369 | 52,116 | |
| Montevideo [130] | Traffic | general | 675 | 690 | 740 | |
| PeMSD7 [15] | Traffic | general | 228 | 19,118 | 1,989 | |
| SZ-taxi [104] | Traffic | general | 156 | 532 | 2,977 | |
| Los-loop [104] | Traffic | general | 207 | 2,833 | 2,017 | |
| Traffic [113] | Traffic | general | 4,438 | 8,996 | 2,160 | |
| HCP [131] | Neuroimaging | general | 360 | * | * | |
| CTDG | Social Evolution [132] | Proximity | general | 74 | 2,099,519 | 565,932 |
| Contact [133] | Proximity | general | 692 | 2,426,279 | 8,065 | |
| Bitcoin-Alpha [134, 135] | Finance | general | 3,783 | 24,186 | 24,186 | |
| Bitcoin-OTC [134, 135] | Finance | general | 5,881 | 35,592 | 35,592 | |
| LastFM [136] | Interaction | bipartite | 1,980 | 1,293,103 | 1,283,614 | |
| MOOC [136] | Interaction | bipartite | 7,144 | 411,749 | 345,600 | |
| Enron [137] | Social | general | 184 | 125,235 | 22,632 | |
| UCI [138] | Social | general | 1,899 | 59,835 | 58,911 | |
| DGraphFin [139] | Social | general | 3,700,550 | 4,300,999 | * | |
| Wikipedia [136] | Social | bipartite | 9,227 | 157,474 | 152,757 | |
| Reddit [136] | Social | bipartite | 10,984 | 672,447 | 669,065 | |
| Taobao [140, 141] | E-commerce | bipartite | 82,566 | 77,436 | * | |
| Taobao-Large [140, 141] | E-commerce | bipartite | 1,630,453 | 5,008,745 | * | |
| eBay-Small [142] | E-commerce | bipartite | 38,427 | 384,677 | * | |
| eBay-Large [142] | E-commerce | bipartite | 1,333,594 | 1,119,454 | * | |
| DTDG | UN Vote [143] | Politics | general | 201 | 1,035,742 | 72 |
| US Legislative [144, 145] | Politics | general | 225 | 60,396 | 12 | |
| Canadian Parliament [144] | Politics | general | 734 | 74,478 | 14 | |
| UN Trade [146] | Finance | general | 255 | 507,497 | 32 | |
| Elliptic [147] | Finance | general | 203,769 | 234,355 | 49 | |
| HEP-TH [148, 149] | Citation | general | 27,770 | 352,807 | 3,487 | |
| MAG [150, 151] | Citation | general | 121,751,665 | 1,297,748,926 | 120 | |
| Twitter-Tennis [152] | Social | general | 1,000 | 40,839 | 120 | |
| Autonomous systems [148] | Communication | general | 7,716 | 13,895 | 733 | |
| Flights [153] | Transport | general | 13,169 | 1,927,145 | 122 | |
| ABIDE [154, 155] | Neuroimaging | general | 200-360 | * | * |
VI Datasets, Evaluation Metrics and Applications
VI-A Datasets
Recently, with more and more attention have been paid to dynamic graph research, there are many time-varying datasets proposed to estimate the performance of different graph models. Here, we divide the time-varying graph datasets into three classes that include the Spatiotemporal Graphs (STGs) dataset, Discrete-Time Dynamic Graphs (DTDGs) dataset and Continuous-Time Dynamic Graphs (CTDGs) dataset. We summarize the statistics of these time-varying graph datasets in Table II.
1) Datasets for STGs In the spatiotemporal graphs research field, most of the representative datasets are used in the traffic domain, with one dataset use for neuroimaging. Los-loop [104] includes loop detector data collected from highways in Los Angeles. This dataset provides insights into traffic speed, volume, and flow on major LA highways, sampled typically at 5-minute intervals. PeMSD7 [15] dataset is part of the larger PeMS dataset from the California Department of Transportation (Caltrans) Performance Measurement System (PeMS), focusing on District 7 (Los Angeles) and includes speed, occupancy, and flow data from freeway sensors. It is used for traffic prediction and anomaly detection research. Traffic [113] consists of traffic data collected over a period of three months, with hourly granularity. It captures general traffic data across multiple urban and freeway road networks from 4,438 sensor stations located in the 7th District of California. METR-LA [106] contains data collected from sensors installed on highways in Los Angeles County, capturing traffic speed at various locations at 5-minute intervals. PeMS-Bay[106] is derived from the PeMS database and includes six months of traffic readings data from Bay Area highways. It provides traffic speed information at 5-minute intervals across numerous sensors. SZ-taxi [104] dataset contains GPS trajectories of taxis in Shenzhen, China, collected over several weeks. This dataset provides information on the location, speed, and timestamp of each taxi, offering a rich source of spatial-temporal traffic data in urban environments. Montevideo [130] captures one month of hourly traffic volume, speed, and congestion levels across various road segments of urban traffic data from the city of Montevideo, Uruguay. HCP [131] (Human Connectome Projectis) is a comprehensive collection of neuroimaging data designed to map the structural and functional connectivity of the human brain from over 1,200 healthy adult participants.
2) Datasets for CTDGs Since there are plenty of dynamic graph dataset for the continuous time dynamic graphs scenario, we choose some representative datasets in each domain to introduce them. Social Evolution [113] dataset comprises dynamic interactions between students at the Massachusetts Institute of Technology (MIT) Human Dynamics Lab. Collected through mobile devices, this dataset captures various forms of temporal social interactions, including phone calls, text messages, and physical proximity, over a period of months. Bitcoin-Alpha [134, 135] is a dynamic network of trust relationships between users on the Bitcoin Alpha platform, where users rate one another based on the trustworthiness of cryptocurrency transactions. LastFM [136] contains one month of who-listens-to-which song information, which was timestamped interactions between users and artists on the LastFM music streaming platform, capturing user preferences over time. UCI [138] is an online temporal communication, such as emails, between students from the University of California, Irvine, with edges in the graph corresponding to messages exchanged between users. Taobao [140, 141] collected from the Chinese e-commerce platform Taobao, contains temporal interactions between users and items, representing user behavior in an online marketplace, which links the meaning of the type of action (e.g., clicking, purchasing, or adding to a wishlist).
3) Datasets for DTDGs In the field of discrete-time dynamic graph research, the following datasets have become widely used, allowing researchers to analyze the evolution of relationships, dependencies, and interactions over discrete time intervals. UN Vote [143] captures voting records of countries in the United Nations General Assembly from 1946 onwards, with a node representing a country, and an edge between nodes reflecting similar voting behavior in a given year. Elliptic [147] contains Bitcoin transaction data from the Elliptic platform, with each transaction labeled as either ”licit”(e.g., exchanges, wallet providers, miners, licit services) or ”illicit”(e.g., scams, malware, terrorist organizations, ransomware, Ponzi schemes). Nodes represent Bitcoin addresses, and edges capture payment flow between these addresses in discrete time steps. HEP-TH [148, 149] is a citation network from arXiv’s High Energy Physics Theory section, where nodes represent papers and directed edges indicate citations between them. The dataset is segmented into monthly intervals, capturing the evolving structure of scientific collaboration and knowledge propagation over time. Twitter-Tennis [152] contains interactions on Twitter related to tennis events, where nodes represent Twitter users, and edges represent mentions, replies, or retweets within specific time intervals. Autonomous Systems [148] is a dynamic graph of Autonomous Systems (AS) over the internet, where nodes represent AS entities and edges indicate connections between them, captured at regular intervals. Flights [153] captures daily flight connections between airports, with nodes representing airports and directed edges indicating flights between them on a specific day. ABIDE [154, 155] (Autism Brain Imaging Data Exchange dataset) consists of functional MRI (fMRI) data from individuals diagnosed with autism spectrum disorder and neurotypical controls, which is often used to create brain connectivity networks over time, where nodes represent brain regions and edges represent functional connectivity.
VI-B Applications
On the application side, recent advancements in TV-GSP techniques promote the development of several fields, which include social science, biological analysis, transportation, environment, and finance due to the outstanding ability of dynamic graph representation in terms of capturing the dynamic relationship between irregularly distributed signals [156, 4, 5, 84].
1) Social science: In social science, time-varying graph signal processing (GSP) has emerged as a powerful tool for analyzing dynamic graph datasets, offering insights into evolving relationships, behaviors, and social interaction patterns. Applications in this domain span various areas, such as opinion dynamics, community detection, anomaly detection, and modeling social behavior.
Opinion Dynamics and Influence Propagation: TV-GSP is employed to study opinion dynamics and influence propagation across social networks, where individual nodes represent users or entities and edges represent their interactions over time. By treating these interactions as signals propagating across a dynamic network, TV-GSP enables the tracking of how opinions, behaviors, or sentiments spread and evolve. Enron [137] dataset serves as a prime example of how opinion and information diffusion can be tracked within an organization by analyzing dynamic interactions over time. This dataset captures email exchanges, offering a temporal sequence of internal communication and opinion exchanges among employees. Additionally, dynamic embedding techniques, such as Reddit and Wikipedia in [136] model user and entity behavior trajectories, allow predictions of future interaction and influence trends. By applying TV-GSP to these datasets, researchers can model the spread of information and identify patterns of influence within the corporate environment.
Community Detection: Community detection in dynamic networks benefits greatly from TV-GSP methodologies, which allow for the identification and tracking of evolving communities over time. In many social networks, communities are not static, and TV-GSP enables the monitoring of their changes, mergers, or splits. UCI in [138] provides an example by analyzing user engagement and behavioral patterns in an online setting. Similarly, [133] offers detailed, time-varying social interaction data that facilitates community evolution tracking and behavior analysis in physical and digital social contexts. This type of analysis, combined with TV-GSP, enables researchers to detect emerging communities and shifting user preferences over time, providing valuable insight into online social interactions.
Detection of Anomalous Patterns: TV-GSP is highly useful for anomaly detection in social networks, where unusual patterns in user behavior or interactions can be flagged in real-time. Such capabilities are crucial for monitoring social media for events like coordinated misinformation efforts or network breaches. In the financial sector, DGraphFin [139] dataset supports such TV-GSP applications, where anomaly detection algorithms are used to detect fraud and irregular transaction patterns. In social contexts, [152] introduces a centrality measure designed for continuously updating network streams, helping identify influential nodes and interactions as new data arrives. [157] employed a multi-level anomaly detection algorithm in dynamic graph settings, effectively detecting anomalies at different graph levels (e.g., node, subgraph) with interactive visualization. These methods prove valuable for anomaly detection within evolving social graphs where central actors can be continually monitored.
Modeling Social Behavior: Modeling social behavior during crisis events, such as natural disasters, pandemics, or public health emergencies, relies on TV-GSP to understand shifts in public sentiment, communication, and connectivity. [113] study exemplifies how temporal GSP techniques are used to assess the collective ”health state” or well-being of a community by tracking temporal patterns in communication and mobility data during significant events. This approach allows for deeper insights into behavior modeling, capturing the temporal evolution of individual behaviors and social influence mechanisms.
2) Biomedical analysis: Dynamic graph-based methods are particularly well-suited to representing temporally evolving interactions, such as neural connectivity patterns, gene expression, and physiological signals. Biomedical applications of TV-GSP encompass brain connectivity studies, predictive modeling in medical diagnostics, and protein interaction analysis.
Dynamic Functional Brain Connectivity Analysis: Functional connectivity analysis is an increasingly important brain function analysis modality, which regards connections as statistical codependency between the signals of different brain regions. TV-GSP has emerged as a valuable framework for analyzing functional brain connectivity, particularly in studies using time-series data from fMRI or EEG. Yu et al. [158] applied TV-GSP to assess dynamic brain connectivity in patients with schizophrenia by analyzing time-varying functional networks derived from fMRI data. This study highlighted how TV-GSP can identify connectivity alterations associated with mental disorders. Calhoun and Adali [159] further expanded by using TV-GSP techniques to track connectivity states across different brain regions, demonstrating applications in both healthy and clinical populations. Recently, Zhao et al. [155] by integrate particle filtering algorithm into graph neural network, achieves superior performance in brain disorder classification.
Medical Diagnostics Modeling: In the field of medical diagnostics, TV-GSP has enabled predictive modeling using dynamic patient data, offering tools for early diagnosis and outcome prediction. Xie et al. [160] introduced a graph convolutional neural network that predicts intensive care interventions based on temporal multivariate data from ICU records, demonstrating the potential of TV-GSP for real-time decision support in critical care environments. Another work by Li et al. [161] explored time-varying graph networks in EEG analysis to monitor cognitive health and diagnose neurological disorders, showing that dynamic graphs can reveal early-stage signs of brain dysfunction in clinical populations.
Protein Interaction and Structural Dynamics Analysis: TV-GSP provides a robust approach for studying dynamic protein interactions, enabling the modeling of temporal and structural changes within protein-protein interaction (PPI) networks. By applying TV-GSP, researchers can capture the evolution of interactions among proteins, which is essential for understanding functional pathways and responses to various stimuli. Song et al. [162] applied dynamic Bayesian networks to track protein signaling pathways over time, using time-varying graphs to model interactions during cellular processes. Furthermore, Li et al. [163] demonstrated how TV-GSP can track dynamic changes in PPI networks, highlighting structural adaptations in proteins under different environmental conditions. Recent work by Calazans et al. [164] leveraged TV-GSP for protein interaction analysis, using graph learning techniques to identify structural motifs and interaction dynamics in protein complexes.
3) Transportation: In transportation systems, TV-GSP has become an essential tool for analyzing dynamic datasets that capture the complex and evolving relationships between various components, such as vehicles, traffic signals, and road networks. Key applications include traffic flow prediction, congestion management, and route optimization.
Traffic Flow Prediction and Congestion Analysis: Traffic flow prediction is one of the most prominent applications of TV-GSP in transportation, where temporal GSP techniques analyze dynamic relationships between road segments or intersections to predict future traffic patterns. By treating traffic flow as a signal on a dynamic graph of road networks, GSP-based models capture both spatial dependencies (interconnectedness of adjacent road segments) and temporal variations (changes in flow over time). The widely used PeMS datasets from California, such as PeMSD7 [15] and PeMSBay [106], contain rich traffic sensor data, enabling effective applications of GSP for short- and long-term traffic prediction. In particular, [106] utilized GSP-based graph convolutional networks on the PeMS datasets to accurately forecast traffic flows by learning spatial-temporal dependencies across the network.
Transportation Network Analysis: TV-GSP also supports the analysis and optimization of public transportation networks, where interactions between transit nodes (e.g., bus stops, train stations) and connections (routes) change continuously over time. By analyzing passenger demand patterns, GSP methods can optimize transit scheduling and enhance network efficiency. For instance, [104] utilized GSP techniques on the SZ-Taxi dataset, which includes GPS-based movement data from taxis in Shenzhen, China, to analyze passenger demand patterns and recommend efficient resource allocation strategies. [130] further supports this application through the PyTorch Geometric Temporal library, which includes tools for implementing GSP methods that optimize public transit networks through spatial-temporal data analysis. In addition, [153] introduces a dataset from the OpenSky network, capturing real-time aircraft signals, which realizes monitoring the transportation condition in the context of air traffic and enables efficient airspace management.
Route Optimization and Navigation: For real-time route optimization, GSP models use dynamic graphs where edge weights vary with traffic conditions, road closures, or accidents, enabling route recommendations that adapt to evolving traffic patterns. Using datasets like METR-LA [106], which contains real-time traffic speed data from Los Angeles, [106] demonstrated that GSP-enabled models could dynamically adapt routing recommendations based on evolving traffic patterns, thus optimizing travel routes in real-time. Additionally, [113] provides a predictive framework for identifying probable path failures within evolving graphs, supporting route planning that can proactively avoid high-risk paths.
4) Environment: Since environmental data is often represented as dynamic graphs with evolving interactions among spatial or temporal nodes, TV-GSP can serve as a valuable framework for monitoring and modeling natural processes. The application of TV-GSP applied in environmental analysis including air quality forecasting, Meteorological prediction, and environmental hazard detection.
Air Quality Prediction: Due to the air quality is highly related to governments energy-saving and emission-reduction policies formulation and resident’s daily life outdoor activities and physical health, air quality prediction has became an critical research issue that attracts increasing attention. Air quality prediction requires the processing of spatially distributed, time-evolving data collected from networks of sensors across urban and rural landscapes. TV-GSP approaches enable the analysis and reconstruction of air pollution signals, providing real-time insights into pollutant levels. For example, [165] introduced a model for reconstructing time-varying graph signals applied to PM2.5 pollution data from California, enabling enhanced tracking of pollution levels over time by leveraging the spatial dependencies between monitoring stations. [166] developed a Sobolev smoothness-based framework that demonstrated effective signal recovery in environmental datasets, highlighting its application to dynamic air quality data. Another work [32] uses spatiotemporal smoothness, facilitating improved predictions of air quality by modeling correlations across sensor networks and adapting to temporal changes.
Meteorological Forecasting: TV-GSP also plays a crucial role in analyzing climate data, where dynamic graph models help capture seasonal and spatial variations. Ortega et al. [5] provided an overview of graph signal processing techniques tailored for environmental datasets, demonstrating their application to large-scale climate data with temporally and spatially varying features. Castro-Correa et al. [167] further advanced this area by developing Gegenbauer Graph Neural Networks specifically designed for reconstructing time-varying signals in complex climate data, allowing for the extraction of meaningful trends over extended periods. Moreover, Lou et al. [168] introduced high-order smoothness and adaptive low-rank techniques to handle the inherent complexity of climate signals, addressing challenges in handling missing data while maintaining the accuracy of climate trend predictions. There are also various meteorological prediction scenarios such as wind prediction [169], [170], [171] and temperature prediction [172], [173], illustrates the capability of TV-GSP in practical use.
Environmental Hazard Detection: Monitoring environmental hazards, such as floods, wildfires, and industrial pollution, benefits from the dynamic adaptability of TV-GSP frameworks, which allow for the real-time detection of anomalies and the tracking of evolving hazard patterns. Valdivia et al. [174] utilized wavelet-based TVGSP methods to visualize and process time-varying environmental data, enabling real-time analysis of environmental events such as wildfires. In a more recent study, Sun et al. [175] developed the Time-Varying Graph Convolutional Network (TVGCN) model to predict multivariate spatiotemporal series, particularly useful in forecasting hazard-related data such as rainfall and water levels in flood-prone areas.
5) Financial analysis: The integration of TV-GSP within financial analysis addresses diverse and complex challenges within dynamic financial systems, where interactions between entities such as stocks, transactions, and institutions change over time. TV-GSP approaches empower researchers and analysts to harness the temporal and structural dynamics of financial interactions, enabling real-time insights into evolving networks. Financial applications of TV-GSP include fraud detection, risk assessment, and stock price forecasting.
Fraud Detection and Anomaly Detection: TV-GSP enables anomaly detection in financial networks, which is critical for identifying irregular transaction patterns and preventing fraud. [135] provides a framework for detecting fraudulent users through the temporal analysis of user ratings, focusing on anomalies in user behaviors and their interactions within the network. Similarly, [147] leverages graph neural networks on a temporal dataset of Bitcoin transactions to detect and prevent money laundering activities in digital finance, leveraging the Bitcoin and Elliptic dataset to identify laundering activities across time-evolving Bitcoin networks.
Risk Assessment and Market Volatility Analysis: Predicting stock market volatility often requires monitoring dynamic, interdependent relationships across different stocks or sectors. The temporal dynamics captured through TV-GSP provide insights into correlations and dependencies that influence market risk and volatility. For instance, [176] employs a time-varying graph model to analyze the dynamic correlations in global stock indices, enabling enhanced forecasting capabilities for market shifts and volatility. In addition, [134] provides a method for predicting future edge weights in financial networks by analyzing both positive and negative relationships, such as cooperative and competitive interactions, within the market. These edge weight prediction techniques assist in understanding market volatility by focusing on how relationship strengths evolve, offering predictive insights into potential price fluctuations.
Stock Price Prediction: Stock price prediction is an essential component of financial trading, enabling traders to make informed decisions regarding the purchasing, selling, and retention of equities. Accurate predictions of future stock prices can help traders optimize their trading strategies and maximize their profits. By applying TV-GSP to financial networks, where nodes represent assets and edges represent correlations, these methods can capture and forecast changes in market conditions, which assists in stock price prediction and correlation analysis. Some common ways of building the TV-GSP for stock price prediction is to combine the graph neural networks (e.g., GCN, GAT) with time series predicting methods (e.g., RNN, LSTM) [177] [178]. While [146] examines the global agricultural trade network, by assessing how countries allocate cropland to domestic versus foreign demand, and considering multiple metrics, presenting a temporal graph of trade dependencies that evolve over time, thus enhancing decisionmaking surrounding trade.
VI-C Evaluation Metrics
1) Evaluation Metrics for Node Level Tasks: Different from static graph signal processing, in the time-varying graph, the most popular task is node-level regression rather than classification, as the core in TV-GSP is the modeling to graph temporal association. Thus, learning from the past to make accurate predictions of the future is the property that most TV-GSP models desire. The metrics for regression are usually drawn from multi-variate time series analysis, which mainly include Mean-Square Error (MSE), Mean Absolute Error (MAE) [118], root Normalized Mean-Square Error(rNMSE) [179], Root Mean Squared Error (RMSE) [180] and Mean Absolute Percentage Error (MAPE) [181]. Taking step ahead prediction as an example, these metrics are defined as functions of node signal as:
In the offline GSP algorithms, they also get to work by replacing the L2 norm with the Frobenius norm, and the functions are then built between node signal matrix and estimated node signal matrix . In addition to regression, there are also some node-level classification tasks on time-varying graphs that usually use Accuracy, Precision, Recall, Area Under Curve for binary classification, and Micro-F1 and macro-f1 for multi-class classification [122, 118, 182, 183]. The formulation of these metrics is the same as seen in static graph signal classification and conventional machine learning tasks.
2) Evaluation Metrics for Edge and Graph Level Tasks: Besides node-level tasks, there are also some edge-level and graph-level tasks on time-varying graphs (e.g., edge prediction and graph classification). Note that, in an unweighted graph, the edge prediction can be viewed as edge classification (i.e., to judge whether the edge exists in the future or not). Thus, the metrics that are popular in classification can also be utilized for them [184, 155, 185]. Besides, for graph topology regression, the regression error of estimated graph topology can be computed by comparing the graph adjacency matrix. In this situation, the input to the regression metrics functions defined in Eq. 31 becomes the adjacency matrix, and the L2 norm becomes Fronbenius norm [186].
3) Evaluation Metrics for Graph Lifelong Learning: In graph lifelong learning, the evaluation metrics change slightly, as its goal is to maintain the ability to make predictions on previous nodes/classes. In this way, the metrics that are usually used include “AP” and “AF” which represent average performance (e.g., the average classification accuracy or regression error) and average forgetting rate. They can be computed by:
where denotes the accuracy of task ; represents the accuracy of task (or classes) after the model being trained on task ; is the total number of tasks; represents the forgetting rate.
VII Conclusion and Future Directions
Time-varying graph signal processing represents a significant advancement in the analysis of dynamic systems. By extending traditional GSP techniques to account for temporal variations in graph structure and node signals, TV-GSP provides a powerful framework for analyzing a wide range of real-world systems. As the field continues to evolve, the integration of TV-GSP with topological signal processing, spatiotemporal graph neural networks, and other advanced techniques will open up new possibilities for research and application.
The field of TV-GSP is still in its infant stage, with many open challenges and opportunities for future research. Some key directions include:
Scalability: As the size of graphs and the length of time series data increase, scalable algorithms that can handle large-scale time-varying graphs are needed. This requires developing efficient computational methods that reduce the complexity of TV-GSP without sacrificing accuracy.
Real-Time Processing: Many applications require real-time analysis of time-varying graphs, such as in cybersecurity or financial markets. Future research should focus on developing real-time TV-GSP methods that can adapt quickly to changes in the graph structure and node signals.
Integration with Machine Learning: There is significant potential in integrating TV-GSP with machine learning techniques, particularly in the context of deep learning. Developing hybrid models that combine the interpretability of GSP with the predictive power of machine learning could lead to new insights and applications.
Theoretical Foundations: Theoretical advancements are needed to better understand the behavior of TV-GSP techniques, especially in terms of stability, convergence, and robustness. Establishing a solid theoretical framework will help guide the development of new methods and applications.
Applications: Finally, exploring new applications of TV-GSP in areas such as neuroscience, epidemiology, and smart cities will drive the field forward. These applications can benefit from the ability to analyze complex, time-varying data on graphs. Conclusion